OpenStudy (anonymous):

Solve the system of equations. 5x - y = 3 and -10x + 2y = -6 (Points : 1) (0, -3) (-3, 0) (3, 12) no solution infinitely many solutions

OpenStudy (anonymous):

5x - y = 3...................(1) -10x + 2y = -6...........(2) multiply (1) by -2: -10x + 2y = -6 ........(3) add (2) and (3): 4y = -12 y = -3 plug y=-3 in (1) 5x -(-3) = 3 5x = 0 x = 0 solution (0,-3)

OpenStudy (unklerhaukus):

infinitely many solutions

OpenStudy (unklerhaukus):

divide the second equation by -2 and the equations are the same